# Online Intrinsic Rewards for Decision Making Agents from Large Language Model Feedback

### v2

@@ A.2 Hyperparameters
For ONI-classification and ONI-ranking, we train the reward model using the Adam optimizer Kingma (2014) with batch size $256$. ONI-classification is trained with learning rate $0.0001$, classification threshold $\eta=0.7$, $\beta=0.1$ for the Score task and $\beta=0.4$ for the others. ONI-ranking is trained with $0.00001$, $\beta=0.05$ and $\nu_{\mathcal{N}}=1.96$ (97.5-th quantile of the standard normal distribution). ONI-retrieval does not train a reward model, and we use $\beta=0.1$ for the Score task and $\beta=0.5$ for the others.
Table A.1 shows the APPO hyperparameters which are common to all experiments and Table A.2 includes the hyperparameters for the online LLM annotation.
@@
Table A.1: Common APPO hyperparameters across all experiments.
@@
Table A.2: LLM annotation hyperparameters.
@@ Appendix B Prompts
Provide a comparative analysis based on first princicples.
@@
Appendix C Additional Results
C.1 LLaMA-3.1-70B-Instruct vs LLaMA-3.1-8B-Instruct
In Figure C.1, we compare the performance of Motif using two different sized LLMs on Score and Oracle (LLaMA-3.1-8B-Instruct and LLaMA-3.1-70B-Instruct). Interestingly, we do not observe a significant difference between the two. This is in contrast to the previous work of (Klissarov et al., 2023), who found a significant difference between using LLaMA-2-70B-chat and LLaMA-2-7B-chat. This suggest that the smaller 8B model is sufficient for evaluating messages on these tasks, hence we use it in our experiments.
Figure C.1: Performance of Motif using two different LLMs. Curves represent means and shaded regions represent standard errors over $5$ seeds.
C.2 Performance vs. LLM Throughput: Tesla V100 vs Tesla A100 GPU
Figure C.2: Performance of ONI-classification is comparable when the LLM server uses a Tesla A100-80GB or V10-32GB GPU.
C.3 Impact of Intrinsic Reward Coefficient $\beta$
Figure C.3: For ONI-retrieval and ONI-classification, sparse reward tasks favor larger intrinsic reward coefficient $\beta$ while smaller values of $\beta$ lead to better results for the dense reward Score task. For ONI-ranking, we do not observe much difference for the Score task, but the sparse reward tasks still slightly favors larger $\beta$.
C.4 Performance when Reducing LLM Annotations
To investigate the effect of reducing the number of LLM annotations, we subsample the LLM annotated messages before sending them back to the hash table (see Section 3.2). Figure 5.5 shows the performance of ONI-retrieval and ONI-classification on Oracle when the subsampling rate is 0.1 and 0.01. ONI-classification is more robust than ONI-retrieval, where its performance is still on par with the original one even when the subsampling rate reduces to 0.01, whereas ONI-retrieval’s performance drops significantly.
C.5 Additional Ablations
Figure C.4 shows the performance and the number of captions stored in the replay buffer for both schemes. We see that for the deduplicated variant, the number of captions does not grow past a certain point (approximately $70k$ unique captions). For the non-deduplicated variant, the number of captions keeps growing linearly over time. The deduplicated variant fails to learn, which highlights the important effect which the annotation dataset can have. We hypothesize that the deduplicated variant may undersample captions that occur frequently in the agent’s experience, for which it is important to reliably estimate reward. For example, the blank message occurs very frequently during policy learning, but is only included in a small fraction of the pairs sent to the LLM for annotation, since it is sampled with the same probability as the other $\sim 70k$ other captions. In contrast, the duplicated variant samples the blank message and other frequent captions with much higher probability. Still, this remains a simple strategy, and designing more sophisticated sampling mechanisms (for example, that account for the epistemic uncertainty of the reward model) would be an interesting direction for future work.
Figure C.4: Performance and number of captions in the replay buffer for agents trained with and without message deduplication. Keeping the natural distribution of captions, and not deduplicating, is important to enable our ranking-based method to learn.
C.6 Goal String and Prompts used in Section 5.2
@@ C.5 Additional Ablations
C.6.1 Prompt used for ELLM-BoW
@@
C.6.2 Prompt used for ONI-retrieval
@@
C.7 Effect of the Episodic Term for ELLM-BoW
Our implementation of ELLM-BoW has included the episodic-count based normalization (7), which is key to the performance of ELLM-BoW. Figure C.5 shows that directly using ELLM-BoW as in Du et al. (2023a) ELLM-BoW, without the episodic term, failed to make progress in all three tasks.
Figure C.5: ELLM-BoW performs well when the intrinsic reward is normalized by an episodic-count based term as in (7). Without it, the success rate is zero for all the three tasks.

### v3

@@ A.2 Hyperparameters
For ONI-classification and ONI-ranking, we train the reward model using the Adam optimizer Kingma (2014) with batch size $256$. ONI-classification is trained with learning rate $0.0001$, classification threshold $\eta=0.5$, $\beta=0.1$ for the Score task and $\beta=0.4$ for the others. ONI-ranking is trained with $0.00001$, $\beta=0.05$ and $\nu_{\mathcal{N}}=0$ (mean of the standard normal distribution). ONI-retrieval does not train a reward model, and we use $\beta=0.1$ for the Score task and $\beta=0.5$ for the others.
Table 1 shows the APPO hyperparameters which are common to all experiments and Table 2 includes the hyperparameters for the online LLM annotation.
@@
Table 1: Common APPO hyperparameters across all experiments.
@@
Table 2: LLM annotation hyperparameters.
@@ Appendix B Prompts
Provide a comparative analysis based on first principles.
@@
B.1 Goal String and Prompts used in Section 5.2
@@
B.1.1 Prompt used for ELLM-BoW
@@
B.1.2 Prompt used for ONI-retrieval
@@
Appendix C Additional Results
C.1 LLaMA-3.1-70B-Instruct vs. LLaMA-3.1-8B-Instruct
In Figure 7, we compare the performance of Motif using two different sized LLMs on Score and Oracle (LLaMA-3.1-8B-Instruct and LLaMA-3.1-70B-Instruct). Interestingly, we do not observe a significant difference between the two. This is in contrast to the previous work of (Klissarov et al., 2023), who found a significant difference between using LLaMA-2-70B-chat and LLaMA-2-7B-chat. This suggest that the smaller 8B model is sufficient for evaluating messages on these tasks, hence we use it in our experiments.
Figure 7: Performance of Motif using two different LLMs. Curves represent means and shaded regions represent standard errors over $5$ seeds.
C.2 Effect of the Episodic Term for ELLM-BoW
Our implementation of ELLM-BoW has included the episodic-count based normalization (6), which is key to the performance of ELLM-BoW. Figure 8 shows that directly using ELLM-BoW as in Du et al. (2023a) ELLM-BoW, without the episodic term, failed to make progress in all three tasks.
Figure 8: ELLM-BoW performs well when the intrinsic reward is normalized by an episodic-count based term as in (6). Without it, the success rate is zero for all the three tasks.
C.3 Additional Ablations
Figure 9: Performance of ONI-classification is comparable when the LLM server uses a Tesla A100-80GB or V100-32GB GPU.
Figure 10: For ONI-retrieval and ONI-classification, sparse reward tasks favor larger intrinsic reward coefficient $\beta$ while smaller values of $\beta$ lead to better results for the dense reward Score task. For ONI-ranking, we do not observe much difference for the Score task, but the sparse reward tasks still slightly favors larger $\beta$.
Impact of Intrinsic Reward Coefficient $\beta$ For ONI-classification and ONI-retrieval, we have used different values of $\beta$ for the Score task and other sparse reward tasks in Section 5.1. Throughout our experiments, we have found that for these two methods, larger values of $\beta$ lead to better performance for the sparse reward tasks, while smaller values of $\beta$ better balance between intrinsic and extrinsic rewards for the dense reward Score task. In comparison, ONI-ranking is relatively robust to this choice, where larger values of $\beta$ are still slightly favored for the sparse reward tasks.
ONI-classification: the impact of the classification threshold As described in Section 3.2, ONI-classification predicts binary rewards by modeling $P(y_{t}=1|o_{t})$ and then thresholding with $\eta$. Instead of using binary reward, an alternative design choice is to use real valued reward
$r^{\mathrm{int}}(o_{t})=P(y_{t}=1|o_{t}),$ (7)
where $r^{\mathrm{int}}(o_{t})\in[0,1]$ is the output of the reward classifier before thresholding. Figure 11 shows that the performance of ONI-classification on the four NetHack tasks is relatively robust to this hyperparameter $\eta$, and using $P(y_{t}=1|o_{t})$ as the reward leads to similar performance. As the training progresses, our reward model outputs values close to 0 or 1777To increase the diversity of the captions in the training dataset, we do not requery the LLM server if a caption is already annotated before. Therefore, every caption in our dataset only has a single label, resulting in this phenomenon., and we hypothesize this is the reason why the final performance remains comparable. Moreover, the natural classification threshold $\eta=0.5$ marginally outperforms the other values in the reward-free setting.
Figure 11: (Top) Performance of ONI-classification is robust to different choices of the classification threshold $\eta$ on the four NetHack tasks. (Bottom) $\eta=0.5$ marginally outperform the other values for training intrinsic-reward-only agents. The result when using $P(y_{t}=1|o_{t})$ as reward (7) is marked with legend oni-classification (no eta).
Figure 12 shows the performance and the number of captions stored in the replay buffer for both schemes. We see that for the deduplicated variant, the number of captions does not grow past a certain point (approximately $70k$ unique captions). For the non-deduplicated variant, the number of captions keeps growing linearly over time. The deduplicated variant fails to learn, which highlights the important effect which the annotation dataset can have. We hypothesize that the deduplicated variant may undersample captions that occur frequently in the agent’s experience, for which it is important to reliably estimate reward. For example, the blank message occurs very frequently during policy learning, but is only included in a small fraction of the pairs sent to the LLM for annotation, since it is sampled with the same probability as the other $\sim 70k$ other captions. In contrast, the duplicated variant samples the blank message and other frequent captions with much higher probability. Still, this remains a simple strategy, and designing more sophisticated sampling mechanisms (for example, that account for the epistemic uncertainty of the reward model) would be an interesting direction for future work.
Figure 12: Performance and number of captions in the replay buffer for agents trained with and without message deduplication. Keeping the natural distribution of captions, and not deduplicating, is important to enable our ranking-based method to learn.